# Snail-Radar: A large-scale diverse dataset for the evaluation of 4D-radar-based SLAM systems

### v1

@@ Abstract
4D radars are increasingly favored for odometry and mapping of autonomous systems due to their robustness in harsh weather and dynamic environments. Existing datasets, however, often cover limited areas and are typically captured using a single platform. To address this gap, we present a diverse large-scale dataset specifically designed for 4D radar-based localization and mapping. This dataset was gathered using three different platforms: a handheld device, an e-bike, and an SUV, under a variety of environmental conditions, including clear days, nighttime, and heavy rain. The data collection occurred from September 2023 to February 2024, encompassing diverse settings such as roads in a vegetated campus and tunnels on highways. Each route was traversed multiple times to facilitate place recognition evaluations. The sensor suite included a 3D lidar, 4D radars, stereo cameras, consumer-grade IMUs, and a GNSS/INS system. Sensor data packets were synchronized to GNSS time using a two-step process: a convex hull algorithm was applied to smooth host time jitter, and then odometry and correlation algorithms were used to correct constant time offsets. Extrinsic calibration between sensors was achieved through manual measurements and subsequent nonlinear optimization. The reference motion for the platforms was generated by registering lidar scans to a terrestrial laser scanner (TLS) point cloud map using a lidar inertial odometry (LIO) method in localization mode. Additionally, a data reversion technique was introduced to enable backward LIO processing. We believe this dataset will boost research in radar-based point cloud registration, odometry, mapping, and place recognition.
@@
4D radar, odometry and mapping, sensor synchronization, radar extrinsic calibration, data reversion, point cloud registration
@@ 1 Introduction
Traditionally, automotive radars are integral to vehicles for obstacle detection and avoidance. In recent years, millimeter-wave radars have also been utilized for odometry and mapping (Harlow et al., 2023) thanks to their robustness to adverse conditions. Two types of radars are commonly used: scanning radars, which rotate to capture 360-view scans (Cen and Newman, 2018), and solid-state single-chip radars, typically with a horizontal field of view of approximately 120 (Cai et al., 2022; Zhuang et al., 2023). Single-chip radars offer higher capture frequencies and are less bulky compared to their scanning counterparts. With the Doppler velocity, a single-chip radar can easily detect moving objects, thus addressing significant challenges encountered in odometry approaches based on cameras (Wang et al., 2021) or lidars (Li et al., 2023). However, conventional automotive radars often have poor vertical resolution, leading to the blending of low and high objects within the same view.
@@
Despite the surge in research utilizing 4D radars for ego-motion estimation, place recognition, and area mapping, existing public 4D radar datasets, e.g., (Kramer et al., 2022; Li et al., 2022; Choi et al., 2023; Zhang, 2023), often fall short in aspects such as accuracy of reference trajectories, diversity in data collection platforms, geographic scope, and repetition of data acquisition on same routes.
@@
Considering the applications in simultaneous localization and mapping (SLAM), we provide sequences with reference trajectories generated using data from a terrestrial laser scanner (TLS). Synchronization based on proven techniques and meticulous calibration ensure the reliability of our data. We believe this dataset will significantly aid in evaluating algorithms for odometry, mapping, and place recognition based on 4D radar point clouds. The dataset and accompanying tools are available at our website 11endnote: 1The dataset website https://snail-radar.github.io/.
@@
    We release a large-scale diverse 4D radar dataset, captured at multiple times over selected routes by three different platforms, encompassing diverse environmental conditions including rainy days and nights, campus roads, and highways. Tools for data conversion between a ROS bag and a folder format, along with calibration results, are also provided.
@@
    A localization pipeline is proposed to generate the reference poses for the provided sequences. The pipeline starts with capturing point clouds of the test area within a campus using a TLS. The stitched point clouds are used as submaps in a LIO method which operates in localization mode. To easily generate reference poses towards the end of a sequence, we also propose a technique to reverse messages for lidar-based odometry.
The structure of the paper is as follows. Section 2 reviews existing radar datasets. The procedure for obtaining 4D radar point clouds is briefly described in Section 3. Details of the dataset, including data collection procedures, sensor setup, and file formats, are provided in Section 4. Ground truth generation, data synchronization, sensor calibration, and known issues are discussed in Sections 5, 6, 7, and 8, respectively. Finally, Section 9 summarizes the paper.
@@ 2 Related work
This section reviews existing radar datasets for odometry and mapping with an emphasis on 4D radars. A comprehensive survey of radar datasets targeting various applications including object detection and tracking have been given in (Zhou et al., 2022).
@@
With spinning radars, several research groups have released radar localization datasets, e.g., the urban Oxford radar robotcar dataset (Barnes et al., 2020), the labeled RADIATE perception dataset (Sheeny et al., 2021), the urban MulRan place recognition (Kim et al., 2020), and the multi-season Boreas dataset (Burnett et al., 2023).
@@
For the recent 4D radars, several datasets have been released. The OdomBeyondVision dataset (Li et al., 2022) was collected by an MAV, a UGV, and a handheld platform in buildings. All three platforms include a FLIR thermal camera and up to three single-chip TI AWR1843 4D radars. The NTU4RadLM dataset (Zhang, 2023) used an Oculii Eagle 4D radar and an iRay thermal camera mounted on a handcart and a car to capture data in clear weather conditions within a campus. The RRxIO dataset (Doer and Trommer, 2021) included a few indoor and outdoor sequences acquired by a thermal camera and a IWR6843 radar mounted on an MAV. The reference trajectories were generated by a motion capture system and a visual inertial SLAM system. The Coloradar dataset (Kramer et al., 2022) was captured by a handheld platform outfitted with a single-chip radar and a cascaded radar. The dataset spans buildings, mine tunnels, and outdoor areas, with reference trajectories generated by a lidar SLAM system. One of its imperfections is that some sequences have wrong Doppler measurements. The USVInland dataset (Cheng et al., 2021) was captured by an unmanned surface vehicles with three TI radars, traveling on inland waterways. The sequences with good GNSS RTK solutions are provided for SLAM evaluations. The MSC RAD4R (Choi et al., 2023) dataset consists of many urban and outskirt sequences captured by a sensor rig mounted on a car. Though the sensor rig includes a GNSS/RTK system and a AHRS system, the RTK solutions often have big closure errors in height and the AHRS system gave wrong headings.
Table 1: Our datasets in comparison to existing radar-based SLAM datasets. Our datasets involves three platforms and diverse scenes, covering small to large geographic scopes.
@@
EU Long-term Conti ARS 308 2D PC + D GPS-RTK/IMU car night, snow urban road
ColoRadar Two TI AWR
@@
USVInland TI AWR1843 3D PC + D GPS-RTK/IMU boat clear waterway
@@
NTU4DRadLM Oculli Eagle 3D PC + D
@@ 3 Radar signal processing background
The digital signal processing, depicted in the bottom of Fig. 1, includes four fundamental components: range processing, Doppler processing, constant false alarm rate (CFAR) detection, and 2D angle of arrival (AoA) processing. Each component is thoroughly illustrated in the TI mmwave SDK documentation. The mathematical principles underlying this pipeline are explained in (Iovescu and Rao, 2017). Here, we provide a high-level functional description of these components, with advanced options omitted for clarity.
@@ 4 Dataset
Table 2: The 44 sequences of our dataset are repeatedly collected at 8 places with three platforms in diverse weather/lighting conditions. The place shorthands bc=basketball court, sl=starlake, ss=software school, if=info faculty, iaf=info and arts faculty, iaef=info, arts, and engineering faculty, st=starlake tower, 81r=August 1 road. The platforms H=handheld, E=ebike, S=SUV. The weather and lighting condition is clear and daytime unless specified. Different to other sequences of the info and arts faculty, the 20240113/2 sequence’s route is anticlockwise.
Place
Setup
@@
More than half of these sequences were captured under special conditions such as rain, dusk, and night. The dataset encompasses various scenes including campus environments, highways, and tunnels. Sample data from camera, lidar, and 4D radar are provided in Table. 3. For place recognition purposes, data for each route were captured at least three different times. Additionally, each sequence was recorded with roughly the same start and end poses.
@@
Figure 2: The routes used for data collection in our dataset. The top plot shows four routes: the basketball court (yellow), the starlake (pink), the software school (brown), and the starlake tower (blue). The bottom plot shows the other four routes: the info arts and engineering faculty (pink), the info and arts faculty (yellow), the August 1 road (blue), and the info faculty (brown).
@@ 4.1 Sensor setups
We assembled the sensors on a aluminum frame designed with CAD, ensuring it was waterproof. The sensor setups for three platforms are drawn in Fig. 3, with sensors for data collection listed in Table 4. The handheld rig consists of a Hesai Pandar XT32 lidar, an Oculii Eagle radar, a ZED2i stereo camera, and a Bynav X36D GNSS/INS system with a single antenna. The rig mounted on the ebike or the SUV includes a Hesai Pandar XT32 lidar, a Continental ARS548 radar, an XSens MTi3DK IMU, an Oculii Eagle radar, a ZED2i stereo camera, and a Bynav X36D GNSS/INS system with dual antennas. Specifically, sequences captured before Nov 1 2023 do not include data from the ARS548 radar and MTi-3DK IMU. These sequences encompass all handheld sequences and some ebike sequences.
@@
A ThinkPad P53 laptop running Ubuntu 20.04 with a 1TB solid-state drive is used for real-time data preprocessing and recording. This laptop supports GPU processing and is compatible with all ROS drivers. It connects to the internet via a WiFi hotspot hosted by a mobile smartphone with 5G networking service. The internet connection is primarily used for updating the computer time via NTP (Network Time Protocol) and receiving RTCM (Radio Technical Commission for Maritime Services) messages used in RTK GNSS positioning, broadcast by the Qianxun FindCM service.
@@ 4.2 File formats
For each sequence, a monolithic rosbag as well as a folder containing separate messages as individual files is provided on OneDrive. The message types in a rosbag and their respective folders are listed in Table 5. In these folders, the point clouds from the Pandar XT32 lidar, the ARS548 radar, and the Eagle radar are saved as pcd files for easy visualization. The compressed images of the ZED2i stereo camera are saved as jpg images. The other topics, e.g., GNSS/INS solutions, and IMU data are saved in txt files.
The ARS548 point clouds are decoded from the detection lists in UDP packets. The X36D messages are also decoded from the UDP packets following the Bynav protocol 22endnote: 2Bynav data communication interface protocol UG016 https://www.bynav.com/cn/resource/support?cat=6.
@@
/radar_trk PointCloud eagleg7/trk bin
@@
For each sequence, we provide reference trajectories of the lidar XT32 frame in the TLS map frame at 10 Hz. However, for large SUV sequences, we only provide the reference poses during the starting subsequence and end subsequence which are covered by the TLS point clouds. To be complete, we provide all the TLS point clouds and their refined poses.
@@
The sensor calibration data are provided in scripts 33endnote: 3Dataset tools https://github.com/snail-radar/dataset_tools.
5 Ground truth
@@
The GNSS RTK/INS solutions from the Bynav system often exhibit height jumps of up to a few meters (Toth and Jóźków, 2016). Therefore, we generate the reference poses using precise TLS point clouds, similar to (Ramezani et al., 2020). But instead of frame-wise ICP, we sequentially align the undistorted lidar frames to the TLS map using a LIO method.
@@
A total of 93 TLS scans were captured by a Leica RTC360 scanner on a sunny day, covering the starlake and the starlake tower routes. In capturing these scans, we ensured sufficient overlap between consecutive scans, with a mean distance of 28.2 m between consecutive scans. These scans were first processed using the Cyclone Register 360 program, with manually identified additional loop constraints between some scans. The TLS scan registration results were further refined by point-to-plane ICP in Open3D and regularized by SE(3) pose graph optimization with uniform weights while considering the two loops within these scans. The pairwise registration results were inspected by three individuals over approximately 10 rounds to ensure proper pairwise matching. The position accuracy of the TLS pairwise registration is expected to be within 5 cm, as implied by the quality check in Fig. 4. The final TLS map was created by stitching together these 93 scans.
@@
Next, for each sequence entirely covered by the TLS map, we localized the lidar scans relative to the TLS map in two steps: first getting the initial pose, and then running LIO in localization mode. To get the initial pose, we selected a lidar scan at the stationary beginning, and registered the scan to a nearby TLS scan in the CloudCompare program by manually picking five pairs of correspondences. The labor is acceptable as our routes used fewer than five unique starting poses separated by a notable distance. Every initial pose was further refined with the classic point-to-plane ICP in Open3D and manually checked in terms of the number of matches and inlier RMSE (root mean squared error). With these initial poses, we run an extended FAST-LIO2 method (Xu et al., 2022) in localization mode. This method loads the constant TLS map and predicts the pose with IMU data, undistorts the lidar scan, and iteratively refines the pose with matches between scan points and map surfels. The LIO method operates in localization mode as the incremental submap building is replaced by loading segments of the TLS map in a sliding manner. Also, to remove the real-time operation constraint and execution randomness, the method was converted to run sequentially in offline mode. The IMU data used in the localization method is from the MTi3DK or the X36D when the MTi3DK is unavailable.
@@
For large-scale sequences, since the TLS map only covers the beginning and end parts of the sequence, we generate the reference trajectory only for the beginning and end subsequences within the TLS coverage, as done in TUM-VI (Schubert et al., 2018). We select data from the first four minutes as the start subsequence and data from the last four minutes as the end subsequence. The start subsequence is processed by the LIO in localization mode which is terminated when the platform’s minimum distance to the TLS trajectories exceeds 8 m. The TLS trajectories are simply the reference poses of multiple medium-scale sequences which cover the entire TLS area.

### v3

# SNAIL Radar: A large-scale diverse benchmark for evaluating 4D-radar-based SLAM

@@ Abstract
4D radars are increasingly favored for odometry and mapping of autonomous systems due to their robustness in harsh weather and dynamic environments. Existing datasets, however, often cover limited areas and are typically captured using a single platform. To address this gap, we present a diverse large-scale dataset specifically designed for 4D radar-based localization and mapping. This dataset was gathered using three different platforms: a handheld device, an e-bike, and an SUV, under a variety of environmental conditions, including clear days, nighttime, and heavy rain. The data collection occurred from September 2023 to February 2024, encompassing diverse settings such as roads in a vegetated campus and tunnels on highways. Each route was traversed multiple times to facilitate place recognition evaluations. The sensor suite included a 3D lidar, 4D radars, stereo cameras, consumer-grade IMUs, and a GNSS/INS system. Sensor data packets were synchronized to GNSS time using a two-step process including a convex-hull-based smoothing and a correlation-based correction. The reference motion for the platforms was generated by registering lidar scans to a terrestrial laser scanner (TLS) point cloud map by a lidar inertial sequential localizer which supports forward and backward processing. The backward pass enables detailed quantitative and qualitative assessments of reference motion accuracy. To demonstrate the dataset’s utility, we evaluated several state-of-the-art radar-based odometry and place recognition methods, indicating existing challenges in radar-based SLAM.
@@
4D radar, odometry and mapping, sensor synchronization, radar extrinsic calibration, data inversion, point cloud registration
@@ 1 Introduction
Traditionally, automotive radars are integral to vehicles for obstacle detection and avoidance (Caesar et al., 2020). In recent years, millimeter-wave radars have also been utilized for odometry and mapping (Harlow et al., 2023) thanks to their robustness to adverse conditions. Two types of radars are commonly used: scanning radars, which rotate to capture 360-view scans (Cen and Newman, 2018), and solid-state single-chip radars, typically with a horizontal field of view of approximately 120 (Cai et al., 2022; Zhuang et al., 2023). Single-chip radars offer higher capture frequencies and are less bulky compared to their scanning counterparts. With the Doppler velocity, a single-chip radar can easily detect moving objects, thus addressing significant challenges encountered in odometry approaches based on cameras (Wang et al., 2021) or lidars (Li et al., 2023). However, conventional automotive radars often have poor vertical resolution, leading to the blending of low and high objects within the same view.
@@
Several datasets for single-chip radars have been released. Among them, the extensive nuScenes dataset features data captured by five 3D (XY+Doppler) radars mounted on an automobile. However, its sequences are limited to discrete 20-second clips with minimal overlap, making it less suitable for SLAM and place recognition tasks. Despite the surge in research utilizing 4D radars for ego-motion estimation, place recognition, and area mapping, existing public 4D radar datasets, e.g., Kramer et al. (2022); Li et al. (2022); Choi et al. (2023); Zhang et al. (2023a), often fall short in aspects such as accuracy of reference trajectories, diversity in data collection platforms, geographic scope, and repetition of data acquisition on same routes.
@@
Considering the applications in simultaneous localization and mapping (SLAM), we provide sequences with reference trajectories generated using data from a terrestrial laser scanner (TLS). Synchronization based on proven techniques and meticulous calibration ensure the reliability of our data. We believe this dataset will significantly aid in evaluating algorithms for odometry, mapping, and place recognition based on 4D radar point clouds. The dataset is made available under the Open Database License 11endnote: 1License: The SNAIL radar dataset is made available under the Open Database License: http://opendatacommons.org/licenses/odbl/1.0/. Any rights in individual contents of the dataset are licensed under the Database Contents License: http://opendatacommons.org/licenses/dbcl/1.0/, at the website 22endnote: 2The dataset website https://snail-radar.github.io/.
@@
    We release a large-scale diverse 4D radar dataset, captured at multiple times over selected routes by three different platforms, encompassing diverse environmental conditions including rainy days and nights, campus roads, and highways. Tools for data loading, visualization, and conversion, along with calibration results, are also provided.
@@
    A sequential localization pipeline is proposed to generate the reference poses for the provided sequences. The stitched TLS point clouds are used as submaps in a lidar-inertial odometry (LIO) method which operates in forward or backward mode. We propose a technique to reverse messages for backward lidar-based odometry. The backward processing enables quantitative accuracy evaluation of the reference trajectory.
  4. 4.
    On the released dataset, we compare three recent radar-based odometry methods, and three recent radar-based place recognition methods, showing the challenges in using 4D radars for SLAM.
The structure of the paper is as follows. Section 2 reviews existing radar datasets. The procedure for obtaining 4D radar point clouds is briefly described in Section 3. Details of the dataset, including data collection procedures, sensor setup, and file formats, are provided in Section 4. Reference trajectory generation, data synchronization, sensor calibration, and known issues are discussed in Sections 5, 6, 7, and 8, respectively. The comparative studies on odometry and place recognition methods are given in Section 9. Finally, Section 10 summarizes the paper.
@@ 2 Related work
This section reviews existing radar datasets for odometry and mapping with an emphasis on 4D radars. A comprehensive survey of radar datasets targeting various applications including object detection and tracking have been given in Zhou et al. (2022).
@@
With spinning radars, several research groups have released radar localization datasets, e.g., the urban Oxford radar robotcar dataset (Barnes et al., 2020), the Oxford offroad radar dataset (Gadd et al., 2024), the labeled RADIATE perception dataset (Sheeny et al., 2021), the urban MulRan place recognition dataset (Kim et al., 2020), and the multi-season Boreas dataset (Burnett et al., 2023).
@@
For the recent 4D radars, several datasets have been released. The OdomBeyondVision dataset (Li et al., 2022) was collected by an MAV, a UGV, and a handheld platform in buildings. All three platforms include a FLIR thermal camera and up to three single-chip TI AWR1843 4D radars. The NTU4RadLM dataset (Zhang et al., 2023a) used an Oculii Eagle 4D radar and an iRay thermal camera mounted on a handcart and a car to capture data in clear weather conditions within a campus. The RRxIO dataset (Doer and Trommer, 2021) included a few indoor and outdoor sequences acquired by a thermal camera and a IWR6843 radar mounted on an MAV. The reference trajectories were generated by a motion capture system and a visual inertial SLAM system. The Coloradar dataset (Kramer et al., 2022) was captured by a handheld platform outfitted with a single-chip radar and a cascaded radar. The dataset spans buildings, mine tunnels, and outdoor areas, with reference trajectories generated by a lidar SLAM system. One of its imperfections is that some sequences have wrong Doppler measurements. The USVInland dataset (Cheng et al., 2021) was captured by an unmanned surface vehicles with three TI radars, traveling on inland waterways. The sequences with good GNSS RTK solutions are provided for SLAM evaluations. The MSC RAD4R (Choi et al., 2023) dataset consists of many urban and outskirt sequences captured by a sensor rig mounted on a car. Though the sensor rig includes a GNSS/RTK system and a AHRS system, the RTK solutions often have big closure errors in height and the AHRS system gave wrong headings.
Table 1: Our datasets in comparison to existing radar-based SLAM datasets. Our dataset involves three platforms and diverse scenes, covering small to large geographic scopes. RA: range azimuth, PC + D: point cloud and Doppler, ADC: analog to digital converter samples.
@@
OORD 2024 Navtech RA GPS SUV night, snow off-road
EU Long-term
2020
Conti ARS 308 2D PC + D GPS-RTK/IMU car night, snow urban road
ColoRadar
2022
Two TI AWR
@@
USVInland
2021
TI AWR1843 3D PC + D GPS-RTK/IMU boat clear waterway
@@
NTU4DRadLM
2023
Oculli Eagle 3D PC + D
@@ 3 Radar signal processing background
The digital signal processing, depicted in the bottom of Fig. 1, includes four fundamental components: range processing, Doppler processing, constant false alarm rate (CFAR) detection, and 2D angle of arrival (AoA) processing. Each component is thoroughly illustrated in the TI mmwave SDK documentation. The mathematical principles underlying this pipeline are explained in Iovescu and Rao (2017). Here, we provide a high-level functional description of these components, with advanced options omitted for clarity.
@@ 4 Dataset
Table 2: The 44 sequences of our dataset are repeatedly collected at 8 routes with three platforms in diverse weather/lighting conditions. The route shorthands bc=basketball court, sl=starlake, ss=software school, if=info faculty, iaf=info and arts faculty, iaef=info, arts, and engineering faculty, st=starlake tower, 81r=August 1 road. The platforms H=handheld, E=ebike, S=SUV. The weather and lighting condition is clear and daytime unless specified. Different to other sequences of the info and arts faculty, the 20240113/2 sequence’s direction is anticlockwise.
Route
&plat.
@@
More than half of these sequences were captured under special conditions such as rain, dusk, and night. The dataset encompasses various scenes including campus environments, highways, and tunnels. Sample data from camera, lidar, and 4D radar are provided in Table 3. For place recognition purposes, data for each route were captured at least three different times. Additionally, each sequence was recorded with roughly the same start and end poses.
@@
Figure 2: The routes used for data collection in our dataset. The top plot shows four routes: the basketball court (yellow), the starlake (pink), the software school (brown), and the starlake tower (blue). The bottom plot shows the other four routes: the info arts and engineering faculty (pink), the info and arts faculty (yellow), the August 1 road (blue), and the info faculty (brown). We slightly offset the paths for clarity.
@@ 4.1 Sensor setups
We assembled the sensors on an aluminum alloy frame designed with CAD, while ensuring it was waterproof. The sensor setups for three platforms are drawn in Fig. 3, with sensors for data collection listed in Table 4. The handheld rig consists of a Hesai Pandar XT32 lidar, an Oculii Eagle radar, a ZED2i stereo camera, and a Bynav X36D GNSS/INS system with a single antenna. The rig mounted on the ebike or the SUV includes the Hesai lidar, the Oculii radar, the ZED2i camera, the Bynav X36D GNSS/INS system with dual antennas, and two additional sensors, a Continental ARS548 radar and an XSens MTi3DK IMU. Specifically, sequences captured before Nov 1 2023 do not include data from the ARS548 radar and MTi3DK IMU. These sequences encompass all handheld sequences and some ebike sequences.
@@
A ThinkPad P53 laptop running Ubuntu 20.04 with a 1TB solid-state drive is used for real-time data preprocessing and recording. This laptop supports GPU processing and is compatible with all used ROS drivers. It connects to the internet via a WiFi hotspot hosted by a mobile smartphone with 5G networking service. The internet connection is primarily used for updating the computer time via NTP (Network Time Protocol) and receiving RTCM (Radio Technical Commission for Maritime Services) messages used in RTK GNSS positioning, broadcast by the Qianxun FindCM service.
@@ 4.2 File formats
The sequences are organized into folders named by their respective dates, with each folder containing all sequences captured on that time of day. For each sequence, both a monolithic zipped ROS bag file and a zipped folder containing individual message files are provided. The dataset is available for download on Microsoft OneDrive and Baidu Netdisk.
The message types in a rosbag and their respective folders are listed in Table 5. In these folders, the point clouds from the Pandar XT32 lidar, the ARS548 radar, and the Eagle radar are saved as pcd files for easy visualization. The compressed images of the ZED2i stereo camera are saved as jpg images. The other topics, e.g., GNSS/INS solutions, and IMU data are saved in txt files. The fields for these sensor messages are detailed in a series of CSV tables available on the dataset website. To facilitate easy frame transformation, we include a ROS1 launch file in each sequence’s folder. This file publishes the static transforms between the available sensors in the sequence.
The ARS548 point clouds are decoded from the detection lists in UDP packets. The X36D messages are also decoded from the UDP packets following the Bynav protocol 33endnote: 3Bynav data communication interface protocol UG016 https://www.bynav.com/cn/resource/support?cat=6.
@@
/radar_trk PointCloud eagleg7/trk pcd
@@
For each sequence, we provide reference trajectories of the lidar XT32 frame in the TLS map frame at 10 Hz. However, for large SUV sequences, we only provide the reference poses during the starting subsequence and end subsequence which are covered by the TLS point clouds. The full trajectories generated by the cascaded pose graph optimization (CPGO) (Huai et al., 2018) are also provided, but their height accuracy is only within 1 meter. These full trajectories are intended for place recognition instead of odometry benchmarking. To be complete, we provide all the TLS point clouds and the refined poses of these TLS scans.
@@
A python-based SDK is also provided to load the sensor data from the sequence folders and to visualize. Also, the SDK supports conversion between the ROS1 bags and sequence folders. The sensor calibration data are provided in MATLAB scripts 44endnote: 4Dataset tools https://github.com/snail-radar/dataset_tools and ROS1 launch files with static_transform_publishers.
5 Reference trajectories
@@
The GNSS RTK/INS solutions from the Bynav system often exhibit height jumps of up to a few meters due to the frequent GNSS outages (Toth and Jóźków, 2016). Therefore, we generate the reference poses using precise TLS point clouds, similar to (Ramezani et al., 2020). But instead of frame-wise ICP, we sequentially align the undistorted lidar frames to the TLS map using a LIO method. A limitation of these reference poses is that they are only available at the beginning and the end of those long sequences. Despite the limitation, proper odometry evaluation is still feasible (Zhang and Scaramuzza, 2018). For completeness, we also provide the full reference trajectories, generated using CPGO.
5.1 TLS-based lidar inertial localization
@@
A total of 93 TLS scans were captured by a Leica RTC360 scanner on a sunny day, covering the starlake and the starlake tower routes. In capturing these scans, we ensured sufficient overlap between consecutive scans, with a mean distance of 28.2 m between consecutive scans. These scans were first processed using the Cyclone Register 360 program, with manually identified additional loop constraints between some scans. The TLS scan registration results were further refined by point-to-plane ICP in Open3D and regularized by SE(3) PGO with uniform weights while considering the two loops within these scans. The pairwise registration results were inspected by three individuals over approximately 10 rounds to ensure proper pairwise matching. The position accuracy of the TLS pairwise registration is expected to be within 5 cm, as implied by the quality check in Table 6. The final TLS map was created by stitching together these 93 scans.
Next, for each sequence, we generate the reference trajectory relative to the TLS map in three steps like the classic Rauch-Tung-Striebel Smoother. First, we determine the initial pose, then perform LIO in both forward and backward localization modes, and finally average the results from forward and backward localization to produce the reference trajectory.
@@
To get the initial pose, we selected a lidar scan at the stationary beginning, and registered the scan to a nearby TLS scan in the CloudCompare program by manually picking five pairs of correspondences. The labor was acceptable as our routes used fewer than five unique starting poses separated by a notable distance. Every initial pose was further refined with the classic point-to-plane ICP in Open3D and manually checked in terms of the number of matches and inlier RMSE (root mean squared error). With these initial poses, we ran an extended FAST-LIO2 method (Xu et al., 2022) in localization mode. This method loads the constant TLS map and predicts the pose with IMU data, undistorts the lidar scan, and iteratively refines the pose with matches between scan points and map surfels. The LIO method operates in localization mode as the incremental submap building is replaced by loading segments of the TLS map in a sliding manner. Also, to remove the real-time operation constraint and execution randomness, the method was converted to run sequentially in offline mode. The IMU data used in the localization method is from the MTi3DK or the X36D when the MTi3DK is unavailable.
@@
For large-scale sequences, since the TLS map only covers the beginning and end parts of the sequence, we ran the localizer only for the beginning and end subsequences within the TLS coverage, as done in TUM-VI (Schubert et al., 2018). We selected data from the first four minutes as the start subsequence and data from the last four minutes as the end subsequence. Each subsequence was processed by the LIO in localization mode which was terminated when the platform’s minimum distance to the TLS trajectories exceeded 8 m. Here the TLS trajectories were simply the reference poses of multiple medium-scale sequences which covered the entire TLS area.